The number of atoms in 9.0g of Aluminium is the same as 8.1g of magnesio, why?
Answer #1

MW of Al = 26.98 g

MW of Mg = 24.305 g

then

ratio ---> 24.305 /26.98 = 0.90085 g of Mg per 1 g of Al

then,, if m = 9 g of Al

0.90085*9 = 8.10765 g

as stated,

1 mol of Al = 1 mol of Mg = 6.022*10^23 atoms

therefore

this is true

since each atom weighs different